twelve could be purchased for 144$.write the two rates (ratios)implied by this statement.how many could be purchased for $1872
Answer by nerdybill(7384)

twelve could be purchased for 144$.write the two rates (ratios)implied by this statement.how many could be purchased for $1872
.
12/144 = x/1872
1/12 = x/1872
1872/12 = x
156 = x (number you can purchase with $1872)